# Aether: Geometric-Aware Unified World Modeling

### v2

@@ Abstract
The integration of geometric reconstruction and generative modeling remains a critical challenge in developing AI systems capable of human-like spatial reasoning. This paper proposes Aether, a unified framework that enables geometry-aware reasoning in world models by jointly optimizing three core capabilities: (1) 4D dynamic reconstruction, (2) action-conditioned video prediction, and (3) goal-conditioned visual planning. Through task-interleaved feature learning, Aether achieves synergistic knowledge sharing across reconstruction, prediction, and planning objectives. Building upon video generation models, our framework demonstrates unprecedented synthetic-to-real generalization despite never observing real-world data during training. Furthermore, our approach achieves zero-shot generalization in both action following and reconstruction tasks, thanks to its intrinsic geometric modeling. Remarkably, even without real-world data, its reconstruction performance is comparable with or even better than that of domain-specific models. Additionally, Aether employs camera trajectories as geometry-informed action spaces, enabling effective action-conditioned prediction and visual planning. We hope our work inspires the community to explore new frontiers in physically-reasonable world modeling and its applications.
Figure 1: An overview of Aether, trained entirely on synthetic data. The figure highlights its three key capabilities: 4D reconstruction, action-conditioned 4D prediction, and visual planning, all demonstrated on unseen real-world data. The 4D reconstruction examples are derived from MovieGen [48] and Veo 2 [62] generated videos, while the action-conditioned prediction uses an observation image from a university classroom. The visual planning example utilizes observation and goal images from an office building. Better viewed when zoomed in. Additional visualizations can be found in our website.
@@ 1 Introduction
The development of visual intelligence systems capable of comprehending and forecasting the physical world remains a cornerstone of AI research. World models have emerged as a foundational paradigm for building autonomous systems that not only perceive but also anticipate environmental dynamics to make reasonable actions. At their core, three capabilities stand out: First, perception equips the system with the ability to capture the intricate four-dimensional (4D) changes—integrating spatial and temporal information—that are essential for understanding the physical world [86, 63, 66, 65, 37, 82]. This continuous sensing of dynamic cues enables a robust representation of the environment. Second, prediction leverages this perceptual information to forecast how the environment will evolve under specific actions, thereby providing a foresight into future states [28, 77, 3, 35, 32, 24, 60]. Finally, planning uses these predictive insights to determine the optimal sequence of actions required to achieve a given goal. Together, these three aspects empower world models to not only represent the current state of the environment but also to anticipate and navigate its future dynamics effectively.
@@ 3.2 Depth Videos Process
Given a depth video $\mathbf{x}_{d}$, we first clip the depth values to a predefined range $[d_{\mathrm{min}},d_{\mathrm{max}}]$. Next, we apply a square root transformation and subsequently compute the reciprocal to convert the depth values into disparity, as described in [57]. Each disparity video clip is then normalized in a scale-invariant manner. Subsequently, the normalized disparity values are linearly mapped from $[0,1]$ to $[-1,1]$. To meet the input requirements of the VAE, the single-channel disparity map is replicated across three channels, as done in prior works [34, 74]. The final depth latent is computed as:
@@
$\hat{\mathbf{x}}_{\mathrm{disp}}$ $=\frac{\mathbf{x}_{\mathrm{disp}}\times 2-1}{\max{(\mathbf{x}_{\mathrm{disp}})}},$ (3)
@@ 3.4 Model Training
As the dataset we use contains video clips with variable lengths and frames per second (FPS), we randomly select $T\in\{17,25,33,41\}$ frames, and the FPS is randomly sampled from $\{8,10,12,15,24\}$. The RoPE [59] coefficients are linearly interpolated to align with them.
@@ 4.1 Zero-Shot Video Depth Estimation
Implementation Details. Video depth estimation is evaluated based on two key aspects: per-frame depth quality and inter-frame depth consistency. These evaluations are performed by aligning the predicted depth maps with the ground truth using a per-sequence scale. We use absolute relative error (Abs Rel) and $\delta<1.25$ (percentage of predicted depths within a $1.25$-factor of true depth) as metrics. For implementation, we adopt the settings outlined in CUT3R [65]. Our baselines include both reconstruction-based methods—such as DUSt3R [66], MASt3R [37], MonST3R [82], Spann3R [63], and CUT3R [65]—and diffusion-based depth estimators, including ChronoDepth [55], DepthCrafter [29], and DepthAnyVideo (DA-V) [74]. It is important to note that when comparing with diffusion-based depth estimators, we apply scale and shift alignment to the ground truth, as most of these methods are not inherently scale-invariant. All videos are resized with original aspect ratios kept to make the short side align with our model’s input size. For videos that exceed the maximum forward processing spatial or temporal size of our model, we employ a sliding window strategy with a stride size of $8$. In regions of overlap between windows, we first estimate a relative scale by calculating the average of element-wise division. This relative scale is then used to adjust the latter window’s depth predictions. Finally, a linspace-weighted average is applied to the overlapping areas, following approaches similar to prior methods [29, 80].
@@ 4.2 Zero-Shot Camera Pose Estimation
Implementation Details. Following MonST3R [82] and CUT3R [65], we evaluate camera pose estimation accuracy on the Sintel [6], TUM Dynamics [58], and ScanNet [10] datasets. Notably, both Sintel and TUM Dynamics contain highly dynamic objects, presenting significant challenges for traditional Structure-from-Motion (SfM) and Simultaneous Localization and Mapping (SLAM) systems. We report Absolute Translation Error (ATE), Relative Translation Error (RPE Trans), and Relative Rotation Error (RPE Rot) after Sim(3) alignment with the ground truth, following the methodology in [65]. The implementation settings are consistent with those used in CUT3R [65]. All videos are resized with original aspect ratios kept and then center cropped to align with our model’s input size. For long videos exceeding our model’s maximum temporal forward processing length, a sliding window strategy with a stride size of $32$ is employed. In overlapping regions between windows, camera poses are aligned following prior methods [64]. Translation alignment is performed using linear interpolation, while quaternion rotations are interpolated with spherical linear interpolation. Additionally, we observed that the generated camera trajectories exhibit noise, likely due to the limited number of denoising steps. To mitigate this, we apply a simple Kalman filter [71] to smooth the trajectories.
@@ Appendix A Author Contributions
Appendix B Raymap to Camera Parameters Algorithm
We adopt a direct approach to recover camera parameters from raymaps, as shown in Algorithm 1.
@@ Appendix B Raymap to Camera Parameters Algorithm
Appendix C Generation Experiments Details
@@
For the validation set of prediction tasks, we collected 93 in-domain scenes and 43 out-domain scenes, with each scene corresponding to a synthetic video clip. The in-domain scenes are collected from the same synthetic environments used in the training dataset, while the out-domain scenes are sourced from entirely different synthetic environments that are not present in the training data.
@@
For action conditioned prediction tasks, Aether accepts both the first frame as observation image input and the camera trajectory of the video clip as action condition input. To ensure fair comparison, we use GPT-4o to generate detailed text annotations for both the initial and final frames of each video clip. These annotations serve as text prompts for CogVideoX, providing comprehensive camera trajectory descriptions. The prompt template for GPT-4o is designed to:
@@
For planning tasks, we construct a validation set following a similar approach to the prediction tasks, comprising 80 in-domain scenes and 40 out-domain scenes from synthetic environments. For each video clip, we extract the initial and final frames as inputs for both Aether and Aether-no-depth models.
@@
Appendix D Addtional Losses in Stage 2 Training
@@
D.1 Multi-Scale Structural Similarity (MS-SSIM) Loss for Color Videos
@@
D.2 Scale- and Shift-Invariant (SSI) Loss for Depth Videos
@@
D.3 Pointmap Loss for Raymaps

### v3

@@ Abstract
The integration of geometric reconstruction and generative modeling remains a critical challenge in developing AI systems capable of human-like spatial reasoning. This paper proposes Aether, a unified framework that enables geometry-aware reasoning in world models by jointly optimizing three core capabilities: (1) 4D dynamic reconstruction, (2) action-conditioned video prediction, and (3) goal-conditioned visual planning. Through task-interleaved feature learning, Aether achieves synergistic knowledge sharing across reconstruction, prediction, and planning objectives. Building upon video generation models, our framework demonstrates zero-shot synthetic-to-real generalization despite never observing real-world data during training. Furthermore, our approach achieves zero-shot generalization in both action following and reconstruction tasks, thanks to its intrinsic geometric modeling. Notably, even without real-world data, its reconstruction performance is comparable with or even better than that of domain-specific models. Additionally, Aether employs camera trajectories as geometry-informed action spaces, enabling effective action-conditioned prediction and visual planning. We hope our work inspires the community to explore new frontiers in physically-reasonable world modeling and its applications.
Figure 1: An overview of Aether, trained entirely on synthetic data. The figure highlights its three key capabilities: 4D reconstruction, action-conditioned 4D prediction, and visual planning, all demonstrated on unseen real-world data. The 4D reconstruction examples are derived from MovieGen [48] and Veo 2 [62] generated videos, while the action-conditioned prediction uses an observation image from a university classroom. The visual planning example utilizes observation and goal images from an office building. Better viewed when zoomed in. Additional visualizations can be found on our website.
@@ 1 Introduction
The development of visual intelligence systems capable of comprehending and forecasting the physical world remains a cornerstone of AI research. World models have emerged as a foundational paradigm for building autonomous systems that not only perceive but also anticipate environmental dynamics to make reasonable actions. At their core, three capabilities stand out: First, perception equips the system with the ability to capture the intricate four-dimensional (4D) changes—integrating spatial and temporal information—that are essential for understanding the physical world [86, 63, 66, 65, 37, 82]. This continuous sensing of dynamic cues enables a geometric representation of the environment. Second, prediction leverages this perceptual information to forecast how the environment will evolve under specific actions, thereby providing a foresight into future states [28, 77, 3, 35, 32, 24, 60]. Finally, planning uses these predictive insights to determine the optimal sequence of actions required to achieve a given goal. Together, these three aspects empower world models to not only represent the current state of the environment but also to anticipate and navigate its future dynamics effectively.
@@ 3.2 Depth Videos Process
Given a depth video $\mathbf{x}_{d}$, we first clip the depth values to a predefined range $[d_{\mathrm{min}},d_{\mathrm{max}}]$. Next, we apply a square root transformation and subsequently compute the reciprocal to convert the depth values into disparity, as described in [57]. Each disparity video clip is then normalized in a scale-invariant manner. Subsequently, the normalized disparity values are linearly mapped from [0,1][0,1][ 0 , 1 ] to [1,1][-1,1][ - 1 , 1 ]. To meet the input requirements of the VAE, the single-channel disparity map is replicated across three channels, as done in prior works [34, 74]. The final depth latent is computed as:
@@
$\hat{\mathbf{x}}_{\mathrm{disp}}$ $=\frac{\mathbf{x}_{\mathrm{disp}}}{\max{(\mathbf{x}_{\mathrm{disp}})}}\times 2-1,$ (3)
@@ 3.4 Model Training
[…]
@@ 4.1 Zero-Shot Video Depth Estimation
Implementation Details. Video depth estimation is evaluated based on two key aspects: per-frame depth quality and inter-frame depth consistency. These evaluations are performed by aligning the predicted depth maps with the ground truth using a per-sequence scale. We use absolute relative error (Abs Rel) and $\delta<1.25$ (percentage of predicted depths within a 1.251.251.25-factor of true depth) as metrics. For implementation, we adopt the settings outlined in CUT3R [65]. Our baselines include both reconstruction-based methods—such as DUSt3R [66], MASt3R [37], MonST3R [82], Spann3R [63], and CUT3R [65]—and diffusion-based depth estimators, including ChronoDepth [55], DepthCrafter [29], and DepthAnyVideo (DA-V) [74]. It is important to note that when comparing with diffusion-based depth estimators, we apply scale and shift alignment to the ground truth, as most of these methods are not inherently scale-invariant. All videos are resized with original aspect ratios kept to make the short side align with our model’s input size. For videos that exceed the maximum forward processing spatial or temporal size of our model, we employ a sliding window strategy with a stride size of 888. In regions of overlap between windows, we first estimate a relative scale by calculating the average of element-wise division. This relative scale is then used to adjust the latter window’s depth predictions. Finally, a linspace-weighted average is applied to the overlapping areas, following approaches similar to prior methods [29, 80].
@@ 4.2 Zero-Shot Camera Pose Estimation
Implementation Details. Following MonST3R [82] and CUT3R [65], we evaluate camera pose estimation accuracy on the Sintel [6], TUM Dynamics [58], and ScanNet [10] datasets. Notably, both Sintel and TUM Dynamics contain highly dynamic objects, presenting significant challenges for traditional Structure-from-Motion (SfM) and Simultaneous Localization and Mapping (SLAM) systems. We report Absolute Translation Error (ATE), Relative Translation Error (RPE Trans), and Relative Rotation Error (RPE Rot) after Sim(3) alignment with the ground truth, following the methodology in [65]. The implementation settings are consistent with those used in CUT3R [65]. All videos are resized with original aspect ratios kept and then center cropped to align with our model’s input size. For long videos exceeding our model’s maximum temporal forward processing length, a sliding window strategy with a stride size of 323232 is employed. In overlapping regions between windows, camera poses are aligned following prior methods [64]. Translation alignment is performed using linear interpolation, while quaternion rotations are interpolated with spherical linear interpolation. Additionally, we observed that the generated camera trajectories exhibit noise, likely due to the limited number of denoising steps. To mitigate this, we apply a simple Kalman filter [71] to smooth the trajectories.
@@ 7 Conclusion and Limitations
Acknowledgments
This work is supported by the National Key R&D Program of China (NO.2022ZD0160102) and Shanghai Artificial Intelligence Laboratory.
@@ Appendix A Author Contributions
Appendix B Robustness of Data Annotation Pipeline
Here we detail three key design choices in our methodology that were specifically implemented to enhance its robustness against common sources of uncertainty in dynamic RGB-D processing.
Robustness in Dynamic Masking Grounding SAM 2 often yields erroneous results for out-of-domain semantic inputs. To enhance the robustness of this process, we select prompts with low uncertainty and discard frames with a high mask-to-image ratio. This approach improves the reliability of our dynamic mask generation, thereby increasing the robustness of all subsequent operations.
Robustness Against Inaccurate Flow Estimation In our video slicing process, we utilize optical flow magnitude and the forward-backward error as key metrics. This approach mitigates the uncertainty inherent in flow estimation during coarse camera pose estimation, leading to more robust initial annotations.
Robustness in Points Trajectory Estimation Similarly, our video slicing is performed based on optical flow magnitude and forward-backward error. In addition, we discard frames with an insufficient number of keypoints. These steps yield a video sequence that is both rich in keypoints and temporally coherent (i.e., without frame discontinuities). Such a sequence is highly conducive to tracking estimation methods and these operations also serve to minimize the uncertainty associated with the tracking process.
Robustness in Failure Sequence Filtering As a final step, we filter out erroneous estimations using three key criteria. We discard an entire sequence if it exhibits an anomalous focal length, if its reprojection error relative to point tracking exceeds a predefined threshold, or if its geometric consistency error surpasses a specified limit.
Conclusion on Overall Robustness Our method consistently yields accurate and clean camera poses with minimal noise. Furthermore, the safeguarding operations detailed above ensure that our processed data is virtually free of failure cases. This outcome is the key to the robustness of our approach.
Appendix C Raymap to Camera Parameters Algorithm
We adopt a direct approach to recover camera parameters from raymaps, as shown in Algorithm 1. For more details, please refer to our GitHub repository.
@@ Appendix C Raymap to Camera Parameters Algorithm
Appendix D Generation Experiments Details
@@
For the validation set of prediction tasks, we collected 93 in-domain scenes and 43 out-of-domain scenes, with each scene corresponding to a synthetic video clip. The in-domain scenes are collected from the same synthetic environments used in the training dataset, while the out-of-domain scenes are sourced from entirely different synthetic environments that are not present in the training data.
@@
For action conditioned prediction tasks, Aether accepts both the first frame as observation image input and the camera trajectory of the video clip as action-conditioned input. To ensure fair comparison, we use GPT-4o to generate detailed text annotations for both the initial and final frames of each video clip. These annotations serve as text prompts for CogVideoX, providing comprehensive camera trajectory descriptions. The prompt template for GPT-4o is designed to:
@@
For planning tasks, we construct a validation set following a similar approach to the prediction tasks, comprising 80 in-domain scenes and 40 out-of-domain scenes from synthetic environments. For each video clip, we extract the initial and final frames as inputs for both Aether and Aether-no-depth models.
@@
Appendix E Additional Losses in Stage 2 Training
@@
E.1 Multi-Scale Structural Similarity (MS-SSIM) Loss for Color Videos
@@
E.2 Scale- and Shift-Invariant (SSI) Loss for Depth Videos
@@
E.3 Pointmap Loss for Raymaps
@@
Figure 5: More reconstruction visualizations.
Figure 6: More visual planning examples.
Figure 7: Qualitative results for ablation study. Please zoom in.
Appendix F More Ablation Study
Acknowledging the importance of ablation studies and working within our computational resources, we conducted a key ablation in Sec. 5.2, where the depth component was removed during training. Results presented in Tab. 5 and 6 demonstrate that excluding the 4D reconstruction target from the multi-task co-training leads to a notable degradation in visual planning performance. This finding strongly supports our paper’s central claim regarding the effective integration of reconstruction and generation within a unified framework. Qualitative results further illustrating this are provided in Fig 7.
Appendix G More analysis in Sec. 4
Our model performs well on the Sintel and Kitti datasets but is comparatively weaker on BONN. The trend is also observed in other diffusion-based methods. We suggest two primary reasons for this. First, BONN’s scene type is indoor. This may be less compatible with the learned priors of video diffusion models. Second, as an older dataset, BONN exhibits lower image quality and contains artifacts such as motion blur. Diffusion models can be particularly sensitive to such image characteristics, potentially impacting their performance.
Appendix H More Training data details
Our synthetic data collection approach directly follows DA-V and TheMatrix, capturing RGB-D videos from AAA games such as Cyberpunk2077 and Horizon5. The initial raw dataset contained about 12.5 million frames. After undergoing camera pose annotation and filtering, this collection was refined to approximately 8.9 million well-annotated frames, which were subsequently used for training. Our camera pose estimation is comparable to other feed-forward methods, which typically trade the higher accuracy of optimization-based techniques for superior run-time efficiency. Reduced performance on ScanNet is likely due to the domain gap from synthetic training data, alongside ScanNet’s imperfect annotations and motion blur.
Appendix I Running time differences.
See Tab. 7.
Table 7: Reconstruction running FPS differences on A100.
Method DUSt3R-GA MASt3R-GA MonST3R-GA Aether (Ours)
Resolution 144 $\times$ 512 144 $\times$ 512 144 $\times$ 512 480 $\times$ 640
FPS 0.76 0.31 0.35 6.14